log (x + 16) = log x + log 16 Solve the logarithmic equation. Be sure to reject any value of x that is not in the domain of the original logarithmic expression. Give the exact answer
Answer by lwsshak3(11628)

log (x + 16) = log x + log 16 Solve the logarithmic equation.
log(x+16)=logx+log16
log(x+16)=log(16x)
x+16=16x
15x=16
x=16/15